OpenStudy (anonymous):

HELP WILL FAN AND MEDAL! Herman is 1/2 as old as his sister in 7 years he will be 2/3 his sister's age. How old is Herman?

OpenStudy (anonymous):

\[{h = 1/2 s, h + 7 = 2/3 (s + 7)}\]h = 7